Solve by the substitution method.
5x+4y=6
x=22-6y

Answer by algebrahouse.com(1659)

5x + 4y = 6
x = 22 - 6y

5(22 - 6y) + 4y = 6 {substituted 22 - 6y in for x into top equation}
110 - 30y + 4y = 6 {used distributive property}
110 - 26y = 6 {combined like terms}
-26y = -104 {subtracted 110 from both sides}
y = 4 {divided both sides by -26}

x = 22 - 6y {bottom original equation}
x = 22 - 6(4) {substituted 22 in, for x, into bottom equation}
x = 22 - 24 {multiplied 4 by 4}
x = -2 {subtracted}
www.algebrahouse.com